OpenStudy (anonymous):

Identify the solution set of |x + 3| + 7 less than or equal to 10 and describe the graph in a complete sentence.

OpenStudy (anonymous):

This is it? \(|x + 3| + 7 \le 10\)

OpenStudy (anonymous):

Okay, first we subtract 7 to both sides, what's 10 - 7?

OpenStudy (anonymous):

3.

OpenStudy (anonymous):

what do u think I agree with @iGreen.

OpenStudy (anonymous):

Yes, so we have: \(|x + 3| \le 3\) Now we break this into two inequalities. \(x + 3 \le 3\) and \(-x - 3 \le 3\) (Note: We multiplied -1 to every term inside the absolute value brackets) Let's solve each of these separately. \(x + 3 \le 3\) Subtract 3 to both sides, what's 3 - 3?

OpenStudy (anonymous):

0.

OpenStudy (anonymous):

Yes, so we have: \(x \le 0\) \(-x - 3 \le 3\) Add 3 to both sides, what's 3 + 3?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Yes, so we have: \(-x \le 6\) Now we divide -1 to both sides, what's 6 / -1?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

Yes, so we have: \(x \ge -6\) (Remember, we switch the sign when multiplying or dividing by a negative number) So we have: \(x \le 0\) and \(x \ge -6\) This will look like: